# FedRecon: Missing Modality Reconstruction in Heterogeneous Distributed Environments

### v2

@@ Appendix B Datasets
Datasets. We employ four multimodal datasets in our experiments, serving two distinct purposes. To evaluate the effectiveness of our multimodal federated learning framework, we use two real-world datasets: (1) MELD (Poria et al., 2018), a multiparty conversational corpus comprising over 9k utterances with both audio and textual transcripts from the TV series Friends, and (2) CrisisMMD (Alam et al., 2018), which contains 18.1k tweets with paired visual and textual information collected during various real-world crisis events. To further assess the generative quality of our MVAE model, we additionally adopt two benchmark multimodal datasets: (3) PolyMNIST (Sutter et al., 2021a), a synthetic dataset featuring five modalities, each constructed by compositing MNIST digits over diverse background images; and (4) CUB (Wah et al., 2011), the Caltech-UCSD Birds dataset, which pairs bird images with fine-grained textual descriptions, posing a challenging setting due to the nuanced modality-specific alignments.
Due to time constraints, we acknowledge certain limitations in our current evaluation. In particular, we did not integrate PolyMNIST and CUB into the federated learning setting, resulting in a gap between our generative evaluation and task-based distributed experiments. In future work, we plan to include these datasets in the federated learning pipeline to establish a more unified benchmark. This extension would enable a three-fold evaluation framework: (1) assessing the overall quality of the generated data across all four datasets, (2) examining the generation performance within a distributed setting to understand how well the model synthesizes data in federated environments, and (3) evaluating the utility of the generated data in downstream tasks to determine whether reconstructions can effectively support task-specific models.
@@
Next, we provide additional details on how the four datasets were utilized in our experiments. Our experimental setup is primarily based on the federated configuration from (Feng et al., 2023). For the MELD dataset, we applied a simplification by retaining only the four most frequent emotion categories: neutral, happy, sad, and angry, in order to reduce class imbalance and improve model convergence. For CrisisMMD, we adopted the original configuration and partitioned the data following the event-based distribution protocol described in prior work. Detailed statistics of these two datasets are provided in Tables 5 and 6.
@@
As for the remaining two datasets, PolyMNIST and CUB, we followed the standard processing protocols established in previous work (Sutter et al., 2021a; Shi et al., 2019; Wah et al., 2011). PolyMNIST is a synthetic dataset constructed from five image modalities. Each sample comprises five distinct images, each representing an MNIST digit placed over a randomly cropped background from five unique texture images—one per modality. All five images share the same digit label, which constitutes the shared information across modalities, while the style of handwriting and the background image vary across modalities, capturing modality-specific noise. The digit labels serve as the class labels in our experiments. CUB (Caltech-UCSD Birds) is a fine-grained dataset that pairs bird images with natural language descriptions. Each datapoint includes an image of a bird and one or more corresponding captions. This dataset poses a more realistic and challenging multimodal learning scenario due to the substantial amount of modality-specific information: the image conveys visual appearance details (e.g., plumage patterns), while the captions vary in focus and granularity, leading to varying degrees of semantic overlap between modalities.
@@
Together, these four datasets allow us to evaluate both the quality and robustness of generative models under varying conditions of modality alignment and shared content. As illustrated in Figures 11, 12, 13, and 14, we present representative examples from the original datasets to better convey the nature and complexity of each modality. These examples help visualize the diversity of the multimodal inputs, as well as the varying levels of semantic overlap across modalities. The images are adapted from the MMVAE+ repository.
Table 5. MELD Statistics. {a, v, t} = audio, visual, text
@@
Table 6. Crisis Dataset Statistics
@@
Figure 11. An example from the MELD dataset, illustrating multimodal data in emotional dialogue analysis.
Figure 12. An example from the CrisisMMD dataset, showing multimodal data used for crisis situation modeling.
Figure 13. An example from the PolyMNIST dataset, illustrating five synthetic modalities where each digit shares the same label but varies in background and handwriting style.
Figure 14. An example from the CUB dataset, showing bird images paired with fine-grained textual descriptions, highlighting the modality-specific nuances in visual and linguistic features.
Appendix C Reproducibility
Regarding the comparative experiments, all our unimodal baselines are based on implementations provided by FedMultimodal. For multimodal comparisons, we reproduced mmFedMC (Yuan et al., 2024), CACMRN (Xiong et al., 2023), and FedMMR (Wang et al., 2024), as none of these methods have publicly available code. In the following, we provide detailed explanations of how each of these baselines was implemented in our setup.
@@ Appendix C Reproducibility
Appendix D Degradation of MVAE
@@
We train our MVAE generator using complete modality data. However, as the modality missing rate increases, the number of complete modality samples decreases, leading to a decline in modality alignment capability. When the missing rate reaches 100%, MVAE degenerates into several unimodal VAEs. In our experimental setup, due to the distributional issues in modality partitioning, reconstruction remains a challenge. Specifically, when a single federated learning client experiences 100% modality missing, reconstruction becomes exceedingly difficult. This is where GGFS (Global Generator Freezing Strategy) comes into play, effectively preventing full degeneration by ensuring stable generative performance even under severe modality loss.
Appendix E Limitations
@@
One limitation of our method lies in the difficulty of achieving stable convergence for the mapping model. Interestingly, we observed conceptual similarities between our design and the training paradigm adopted in LLaVA (Liu et al., 2023), which inspired us to adapt their training methodology. This proved beneficial—particularly for challenging datasets such as CrisisMMD, where we incorporated KL regularization to align the distributions of the two latent variables. This approach partially alleviated convergence issues.
@@ Appendix F Tricks
Appendix G Visulazation
To qualitatively assess the effectiveness of our reconstruction framework, we visualize generated samples obtained by reusing the encoder and decoder from MMVAE+, followed by our trained modality-mapping network. These visualizations demonstrate how well our mapping model can align and reconstruct missing modalities based on available inputs.
In certain scenarios, the mapping achieves noticeably better cross-modal alignment, yielding more coherent and semantically consistent samples. This highlights the model’s ability to preserve modality-specific characteristics while capturing shared semantic content across modalities. Figures 1523 showcases representative examples from different datasets to illustrate the alignment quality and generative fidelity.
Figure 15. Visualization example 1: Unconditional self-reconstruction using the MMVAE+ encoder and decoder.
Figure 16. Visualization example 2: Unconditional self-reconstruction using the MMVAE+ encoder and decoder.
Figure 17. Visualization example 3: Cross-modal generation using our trained mapping network and MMVAE+ backbone.
Figure 18. Visualization example 4: Cross-modal generation using our trained mapping network and MMVAE+ backbone.
Figure 19. Visualization example 5: Cross-modal generation using our trained mapping network and MMVAE+ backbone.
Figure 20. Visualization example 6: Cross-modal generation using our trained mapping network and MMVAE+ backbone.
Figure 21. Visualization example 7: Cross-modal generation using our trained mapping network and MMVAE+ backbone.
Figure 22. Visualization example 8: Cross-modal generation using our trained mapping network and MMVAE+ backbone.
Figure 23. Visualization example 9: Cross-modal generation using our trained mapping network and MMVAE+ backbone.

### v3

@@ Appendix D Datasets
Datasets. We employ four multimodal datasets in our experiments, serving two distinct purposes. To evaluate the effectiveness of our multimodal federated learning framework, we use two real-world datasets: (1) MELD (Poria et al. 2018), a multiparty conversational corpus comprising over 9k utterances with both audio and textual transcripts from the TV series Friends, and (2) CrisisMMD (Alam, Ofli, and Imran 2018), which contains 18.1k tweets with paired visual and textual information collected during various real-world crisis events. To further assess the generative quality of our MVAE model, we additionally adopt two benchmark multimodal datasets: (3) PolyMNIST (Sutter, Daunhawer, and Vogt 2021a), a synthetic dataset featuring five modalities, each constructed by compositing MNIST digits over diverse background images; and (4) CUB (Wah et al. 2011), the Caltech-UCSD Birds dataset, which pairs bird images with fine-grained textual descriptions, posing a challenging setting due to the nuanced modality-specific alignments.
@@
Next, we provide additional details on how the four datasets were utilized in our experiments. Our experimental setup is primarily based on the federated configuration from (Feng et al. 2023). For the MELD dataset, we applied a simplification by retaining only the four most frequent emotion categories: neutral, happy, sad, and angry, in order to reduce class imbalance and improve model convergence. For CrisisMMD, we adopted the original configuration and partitioned the data following the event-based distribution protocol described in prior work. Detailed statistics of these two datasets are provided in Tables 5 and 6.
@@
As for the remaining two datasets, PolyMNIST and CUB, we followed the standard processing protocols established in previous work (Sutter, Daunhawer, and Vogt 2021a; Shi et al. 2019; Wah et al. 2011). PolyMNIST is a synthetic dataset constructed from five image modalities. Each sample comprises five distinct images, each representing an MNIST digit placed over a randomly cropped background from five unique texture images—one per modality. All five images share the same digit label, which constitutes the shared information across modalities, while the style of handwriting and the background image vary across modalities, capturing modality-specific noise. The digit labels serve as the class labels in our experiments. CUB (Caltech-UCSD Birds) is a fine-grained dataset that pairs bird images with natural language descriptions. Each datapoint includes an image of a bird and one or more corresponding captions. This dataset poses a more realistic and challenging multimodal learning scenario due to the substantial amount of modality-specific information: the image conveys visual appearance details (e.g., plumage patterns), while the captions vary in focus and granularity, leading to varying degrees of semantic overlap between modalities.
@@
Together, these four datasets allow us to evaluate both the quality and robustness of generative models under varying conditions of modality alignment and shared content. As illustrated in Figures 8, 9, 10, and 11, we present representative examples from the original datasets to better convey the nature and complexity of each modality. These examples help visualize the diversity of the multimodal inputs, as well as the varying levels of semantic overlap across modalities. The images are adapted from the MMVAE+ repository.
Table 5: MELD Statistics. {a, v, t} = audio, visual, text
@@
Table 6: Crisis Dataset Statistics
@@
Figure 8: An example from the MELD dataset, illustrating multimodal data in emotional dialogue analysis.
Figure 9: An example from the CrisisMMD dataset, showing multimodal data used for crisis situation modeling.
Figure 10: An example from the PolyMNIST dataset, illustrating five synthetic modalities where each digit shares the same label but varies in background and handwriting style.
Figure 11: An example from the CUB dataset, showing bird images paired with fine-grained textual descriptions, highlighting the modality-specific nuances in visual and linguistic features.
Appendix E Implementation
For our MFL framework, we implement FedRecon atop the FedMultimodal benchmark (Feng et al. 2023). Following their protocol, we utilize MobileNetV2 (Howard et al. 2017), MobileBERT (Sun et al. 2020), and MFCC (Mel-Frequency Cepstral Coefficients) (Davis and Mermelstein 1980) to extract features from visual, textual, and audio modalities respectively, enabling efficient and scalable federated training. Therefore, a simple Conv+RNN architecture is used to process visual and audio data, while text is handled by an RNN-only model. For the MELD dataset, which contains only audio and text modalities, we adapt our model structure based on the VAE design provided in a prior implementation [^3] https://github.com/yjlolo/vae-audio to better handle the audio modality. For CrisisMMD, we adopt the vision-language VAE architecture proposed in the original MVAE framework to accommodate the image-text modality pair. We use FedAvg as the baseline method for MELD, while for CrisisMMD we adopt FedOpt. We report Unweighted Average Recall (UAR) on MELD and F1-score on CrisisMMD, along with Top-1 accuracy on both datasets for a more comprehensive evaluation. To evaluate generative baselines, we build upon the official MMVAE+ implementation [^4] https://github.com/epalu/mmvaeplus and refer to their reported metrics for consistency. We also reproduce MMVM [^5] https://github.com/thomassutter/mmvmvae results as an additional reference point. Generative quality is assessed using Fréchet Inception Distance (FID) (Heusel et al. 2017) and generative coherence (Shi et al. 2019) metrics. For the coefficient $\gamma$ controlling the KL regularization term, we applied it in the experiments on the CIRSMMD and CUB datasets. All experiments are repeated with three different random seeds to ensure robustness and are conducted on a single NVIDIA A100 GPU with 80GB memory.
For all task models, we adopt the training hyperparameter settings provided by the FedMultimodal benchmark to ensure fair comparison and reproducibility. We also detail the hyperparameters used for MVAE training, which include 20 local epochs per communication round, a learning rate of $1\times 10^{-3}$, and equal weighting coefficients of 1 for both the KL divergence loss ($L_{\mathrm{KL}}$) and the reconstruction loss ($L_{\mathrm{recon}}$). The mapping model is trained for 40 epochs per communication round with a learning rate of $5\times 10^{-3}$. Through empirical testing, we found these hyperparameter settings provide stable training performance and are well suited across all datasets used in our experiments.
Regarding the comparative experiments, all our unimodal baselines are based on implementations provided by FedMultimodal. For multimodal comparisons, we reused the first three methods—mmFedMC (Yuan et al. 2024), CACMRN (Xiong et al. 2023), and FedMMR (Wang et al. 2024)—as none of them have publicly available code. For CLAP (Cui et al. 2024) and FedMobile (Liu, Wang, and Yuan 2025), we based our implementation on their open-source repositories. In the following, we provide detailed explanations of how each of these baselines was implemented in our setup.
@@ Appendix E Implementation
For CLAP and FedMobile, we adjusted their generator architectures to better fit our data modalities and conducted distributed training under our experimental setup. We categorized them into two types based on whether sequence data (e.g., text, audio) is present or not (e.g., images). For each category, we adopted the corresponding original architecture from their official implementations. Ultimately, we evaluated the performance using our defined task model alongside the trained generators.
We have open-sourced the code for training the mapping models. You can adapt and use any pretrained MVAE model to work with our repository. Our MVAE training is based on MMVAE and MMVAE+, and you may refer to their open-source code for implementation details. Our federated learning testing is based on FedMultimodal, where you can download datasets and initialize models accordingly.
Appendix F More Rusults
In this subsection, we provide additional qualitative results to further illustrate the reconstruction ability and cross-modal alignment of our framework. These visualizations are aimed at demonstrating how our method recovers missing modalities under diverse scenarios and dataset conditions. Figures 1321 show generated results on the CUB and PolyMNIST datasets, where each sample illustrates how the trained modality-mapping network aligns latent representations across modalities.
In addition, Figures 22 and 23 present cross-modal generation results on the CrisisMMD dataset. Specifically, Figure 22 shows text-to-image generation, where the model generates visual content conditioned on input text. Figure 23 shows image-to-text generation, where the model synthesizes textual descriptions based on visual input. These examples demonstrate the model’s ability to perform bidirectional modality completion across vision and language, capturing both modality-specific detail and cross-modal semantic consistency.
Appendix G More Ablation Results
To better understand the contribution of each component in our framework, we provide additional ablation results on the MELD and CrisisMMD datasets. Table 7 reports UAR scores on MELD under different proportions of missing modalities, while Table 8 shows F1-scores on CrisisMMD when $\alpha=0.1$. In both cases, removing either the MMR module or the GGFS module leads to noticeable performance drops, highlighting their complementary roles. Notably, MMR contributes more significantly to the performance gain compared to GGFS.
Table 7: UAR scores on MELD with different missing ratios. M: the cross-modal alignment reconstruction generator (MMR), G: the Global Generator Freezing Strategy (GGFS).
Setting 10% 20% 30% 40% 50% 60% 70% 80% 90%
w/o M + G 67.09 67.05 66.78 66.02 63.46 64.66 64.81 62.51 61.39
w/o G 69.82 68.85 68.62 68.75 66.84 66.40 64.97 63.56 64.27
FedRecon 70.11 69.94 69.65 69.18 67.12 66.53 65.28 63.91 64.81
Table 8: F1 scores on CrisisMMD ($\alpha=0.1$) under different missing ratios.
Setting 10% 20% 30% 40% 50% 60% 70% 80% 90%
w/o M + G 37.53 38.29 38.68 36.06 35.51 32.13 30.33 28.08 15.36
w/o G 41.06 41.97 41.52 39.85 38.01 34.82 32.55 28.37 18.94
FedRecon 44.50 44.48 42.83 40.23 38.83 35.25 35.00 28.82 21.10
Appendix H GGFS Analysis and Visualization
We present the distribution of the number of samples per client in Figure 12, illustrating the scenario hypothesized in the main text. Specifically, some clients hold only a few samples that are difficult to reconstruct using their local generator, due to the presence of many other classes locally, which weakens the local generator’s ability to approximate these hard samples. In contrast, the global generator can better simulate these hard samples, as other clients possess a large number of similar hard samples.
Figure 23 shows results using GGFS, while Figure 24 presents results without GGFS. It can be observed that in the volcanic eruption scenario, the generator without GGFS produces misunderstandings.
Figure 12: Number of samples per class allocated to each client in CrisisMMD with $\alpha=5.0$ (a) and $\alpha=0.1$ (b).
Table 9: Reconstruction of a missing modality on the PolyMNIST dataset using different combinations of available modalities.
Available Modalities Coherence ↑ FID ↓
$m_{0}$ 0.839 76.87
$m_{1}$ 0.817 78.18
$m_{2}$ 0.834 79.20
$m_{3}$ 0.843 78.56
$m_{4}$ 0.845 78.92
$m_{0}+m_{1}$ 0.847 79.29
$m_{3}+m_{4}$ 0.858 78.18
$m_{1}+m_{3}+m_{4}$ 0.852 79.86
$m_{0}+m_{1}+m_{2}+m_{3}+m_{4}$ 0.861 80.21
We further examine whether a single available modality or multiple concatenated modalities should be used to reconstruct a missing one. As shown in Table 9, on the PolyMNIST dataset, reconstruction using any single modality ($m_{0}$ to $m_{4}$) yields comparable performance. While incorporating multiple modalities can slightly improve both coherence and FID, the gain is marginal. Thus, using a single modality is generally sufficient and more efficient in terms of reconstruction cost.
Appendix I Training Cost Analysis
We report the training cost of FedRecon across different datasets, as summarized in Table 10. The total runtime consists of three components: MVAE, the mapping model, and the task model. Results show that the overall training time remains practical even on large-scale or high-resolution datasets.
Table 10: FedRecon runtime cost. Total training time is broken down into MVAE, Mapping Model, and Task Model components.
Model Runtime/Round Runtime/Round Total Time Total Time
(CrisisMMD) (MELD) (PolyMNIST) (CUB)
MVAE 26 s 21 s 189 min 301 min
MMVAE 33 s 34 s 237 min 923 min
MMVAE+ 35 s 33 s 249 min 1036 min
Original Mapping Model 4 s 4 s 37 min 34 min
Simplified Mapping Model 4 s 4 s 9 min 34 min
Task Model 4 s 3 s 15 min 17 min
Total (FedRecon) 34 s 28 s 241 min 352 min
We primarily adopt the original mapping model in all experiments for consistency. However, we also evaluate a simplified variant that removes the dependency on source modality identity. As shown on PolyMNIST, this simplified version can further reduce training time with only marginal performance differences (see Appendix J). In practice, we observe that both the MVAE and the mapping model tend to converge early, making early stopping an effective strategy to save computational cost. Additionally, when using pre-extracted features under the FedMultimodal setup, the peak GPU memory usage stays below 1000MB even on a GTX 1060, enabling fast and memory-efficient training.
Appendix J Simplified Mapping Model Evaluation
We simplify the modality mapping model by removing the dependency on the source modality index. We evaluate this simplified design on PolyMNIST, with results shown in Table 11.
Table 11: Evaluation of simplified mapping design on the PolyMNIST dataset. $T_{n,m}$ denotes a modality-specific mapping model, while $T_{n}$ uses a shared mapping regardless of the source modality.
Model Coherence ↑ FID ↓
$T_{n,m}$ 0.845 78.92
$T_{n}$ 0.827 80.51
Appendix K Extension to MMVAE+ Backbone
We further evaluate FedRecon using MMVAE+ as the backbone encoder, demonstrating its compatibility with more expressive multimodal models. As shown in Figure 25, this setup yields improved coherence and generation quality under various qualitative scenarios.
Appendix L Comparison with CLAP
CLAP is a pioneering work that first leverages MVAE for modality reconstruction under missing-modality settings. However, its missingness pattern differs fundamentally from ours. Specifically, CLAP assumes a “hard” missingness setup, where all samples within a client lack the same modalities, and the remaining modalities are complete and perfectly aligned (e.g., client 1 always lacks modality $X$, while having full access to $Y$ and $Z$). In contrast, FedRecon is designed for a more realistic “soft” missingness scenario, where the presence or absence of each modality varies on a per-sample basis within each client, typically following a Bernoulli distribution.
This distinction is critical: sample-level missingness reflects practical settings such as sensor dropouts or incomplete annotations, which are common in decentralized real-world applications. As discussed in the Introduction, FedRecon is the first to apply MVAE in such soft missing-modality federated settings.
Moreover, CLAP does not provide empirical analysis on why MVAE is preferable to other generative alternatives, especially under distributed constraints. In contrast, we provide experimental evidence to support the use of MVAE as a backbone for modality completion. While our setup is not fully aligned with CLAP’s original assumptions, it still satisfies the client-level connectivity requirements. Nonetheless, due to the high number of clients and the reduced per-client data under high missingness, CLAP’s performance degrades significantly, highlighting its limited applicability in realistic federated conditions.
Appendix M Limitations
@@ Appendix I Training Cost Analysis
Degradation of MVAE. We train our MVAE generator using complete modality data. However, as the modality missing rate increases, the number of complete modality samples decreases, leading to a decline in modality alignment capability. When the missing rate reaches 100%, MVAE degenerates into several unimodal VAEs. In our experimental setup, due to the distributional issues in modality partitioning, reconstruction remains a challenge. Specifically, when a single federated learning client experiences 100% modality missing, reconstruction becomes exceedingly difficult. This is where GGFS comes into play, effectively preventing full degeneration by ensuring stable generative performance even under severe modality loss.
@@
Training Stability. Another limitation of our method lies in the difficulty of achieving stable convergence for the mapping model. Interestingly, we observed conceptual similarities between our design and the training paradigm adopted in LLaVA (Liu et al. 2023), which inspired us to adapt their training methodology. This proved beneficial—particularly for challenging datasets such as CrisisMMD, where we incorporated KL regularization to align the distributions of the two latent variables. This approach partially alleviated convergence issues.
@@
Figure 13: Visualization example 1: Unconditional self-reconstruction using the MVAE encoder and decoder.
Figure 14: Visualization example 2: Unconditional self-reconstruction using the MVAE encoder and decoder.
Figure 15: Visualization example 3: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 16: Visualization example 4: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 17: Visualization example 5: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 18: Visualization example 6: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 19: Visualization example 7: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 20: Visualization example 8: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 21: Visualization example 9: Cross-modal generation using our trained mapping network and MVAE backbone.
Figure 22: Examples of Image generated from Text on the CrisisMMD dataset, using the backbone trained with the MVAE framework.
Figure 23: Examples of Text generated from Image on the CrisisMMD dataset, using the backbone trained with the MVAE framework.
Figure 24: Examples of Text generated from Image on the CrisisMMD dataset, using the backbone trained with the MVAE framework without GGFS.
Figure 25: Visualization example 10: Cross-modal generation using our trained mapping network and MMVAE+ backbone.